Test the following series for convergence


n=1 [✓(n^4 +9) -✓(n^4 -9)]


1
Expert's answer
2021-06-14T18:42:33-0400
"\displaystyle\sum_{n=1}^{\infin}(\sqrt{n^4+9}-\sqrt{n^4-9})"




"=\displaystyle\sum_{n=1}^{\infin}(\sqrt{n^4+9}-\sqrt{n^4-9})\dfrac{\sqrt{n^4+9}+\sqrt{n^4-9}}{\sqrt{n^4+9}+\sqrt{n^4-9}}"


"=\displaystyle\sum_{n=1}^{\infin}\dfrac{n^4+9-n^4+9}{\sqrt{n^4+9}+\sqrt{n^4-9}}"

"=18\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{\sqrt{n^4+9}+\sqrt{n^4-9}}"

Use the Limit Comparison Test with


"a_n=\dfrac{1}{\sqrt{n^4+9}+\sqrt{n^4-9}}, b_n=\dfrac{1}{n^2}"


and obtain


"\lim\limits_{n\to\infin}\dfrac{a_n}{b_n}=\lim\limits_{n\to\infin}\dfrac{\dfrac{1}{\sqrt{n^4+9}+\sqrt{n^4-9}}}{\dfrac{1}{n^2}}"

"=\lim\limits_{n\to\infin}\dfrac{1}{\sqrt{1+\dfrac{9}{n^4}}+\sqrt{1-\dfrac{9}{n^4}}}=\dfrac{1}{2}>0"

Since


"\displaystyle\sum_{n=1}^{\infin}b_n=\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{n^2}"

is convergent ("p" -series with "p=2" ), the series "\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{\sqrt{n^4+9}+\sqrt{n^4-9}}" converges by the Limit Comparison Test.


Therefore the given series "\displaystyle\sum_{n=1}^{\infin}(\sqrt{n^4+9}-\sqrt{n^4-9})" is convergent by the Limit Comparison Test.